OpenStudy (anonymous):

9|9-8x|=2x+3 Please explain the steps to completing. I will medal and fan!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you have to solve two equations i think

OpenStudy (anonymous):

for \(x>\frac{9}{8}\) this is the same as \[9(9-8x)=2x+3\]

OpenStudy (anonymous):

oh maybe i just screwed that up lets make it easiery

OpenStudy (anonymous):

\[|9-8x|=|8x-9|\] so we can start with \[9|8x-9|=2x+3\]

OpenStudy (anonymous):

now if \(x>\frac{9}{8}\) that is the same as \[9(8x-9)=2x+3\] which i guess you can solve

OpenStudy (anonymous):

it is ugly, but you should get \(x=\frac{6}{5}\) i think

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

lets back up and go slow

OpenStudy (anonymous):

is this line \[|9-8x|=|8x-9|\] clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now \(|8x-9|\) is one of two things if \(x> \frac{9}{8}\) then \(8x-9>0\) and if \(8x-9>0\) then \(|8x-9|=8x-9\) because it is positive

OpenStudy (anonymous):

clear or not ?

OpenStudy (anonymous):

clear

OpenStudy (anonymous):

so if \(x>\frac{9}{8}\) you can go ahead and solve \[9(8x-9)=2x+3\]

OpenStudy (anonymous):

that takes a bit of algebra, but i assume it is doable

OpenStudy (anonymous):

if not, let me know

OpenStudy (anonymous):

I don't know how to solve it @satellite73

OpenStudy (anonymous):

multiply out and get \[72x-81=2x+3\]

OpenStudy (anonymous):

Do i combine like terms now?

OpenStudy (anonymous):

subtract \(2x\) get \[70x-81=3\] then add \(81\) to get \[70x=84\] and finally divide by \(70\) to get \[x=\frac{84}{70}=\frac{6}{5}\]

OpenStudy (anonymous):

ooooooohhhhh ok thank you so much again!!!!